# Uncertainty Quantification with the Empirical Neural Tangent Kernel

### v1

@@ Appendix E Experiment Details
E.1 Toy Regression
We use a 1-layer MLP, with a width of $50$ and SiLU activation. For the maximum a posteriori (MAP) network, we train for $10000$ epochs, with a learning rate of $0.001$, using the Adam optimizer and the PyTorch polynomial learning rate scheduler, with parameters total__\__iters $=$ epochs, power $=~{}0.5$. For DE, each network in the ensemble outputs a heteroskedastic variance, and is trained using a Gaussian NLL, with $2000$ epochs and a learning rate of $0.05$. We combine the predictions of the ensembles as per (Lakshminarayanan et al., 2017). Both DE and NUQLS use $10$ realizations. The $\gamma$ hyper-parameter in NUQLS is set to $5$, and each linear realization is trained for $1000$ epochs with a learning rate of $0.001$, using SGD with a momentum parameter of $0.9$. In SWAG, the MAP network is trained for a further $10000$ epochs, using the same learning rate, and the covariance is formed with a rank-$10$ approximation. A prior precision of $0.1$ and $1$ is used for LLA and LA respectively, as well as the full covariance matrix. The variational inference method used is Bayes By Backprop (Blundell et al., 2015), as deployed in the Bayesian Torch package (Krishnan et al., 2022). The prior parameters are ($\mu=0,\sigma=1$), and the posterior is initialized at ($\mu=0,\rho=-3$). For SWAG, LLA, LA and VI, $1000$ MC sample were taken at test time. These design choices gave the best performance for this problem.
E.2 UCI Regression
Table 8: Training procedure for UCI regression results in Table 1 and Table 6.
@@ E.2 UCI Regression
No. experiments $10$ $10$ $10$ $10$ $10$ $10$ $10$ $10$ $3$
We now provide the experimental details for the UCI regression experiments (as seen in Table 1 and Table 6). For each dataset, we ran a number of experiments to get a mean and standard deviation for performance metrics. In each experiment, we took a random $70\%/15\%/15\%$ split of the dataset for training, testing, and validation. The training hyper-parameters for the MAP, DE and NUQLS networks, size of the MLP used, and the number of experiments conducted for each dataset can be found in Table 8.
  • NN: For the PolyLR learning rate scheduler, a PyTorch polynomial learning rate scheduler was used, with parameters total__\__iters=10×10$\times$10 ×epochs, power$=0.5$. The MLP used a $\tanh$ activation, so as to have smooth gradients. MLP weights were initialized as Xavier normal, and bias as standard normal. The dataset was normalized, so that the inputs and the outputs each had zero mean and unit standard deviation.
@@
  • LLA: LLA requires two parameters for regression: a dataset noise parameter, and a prior variance on the parameters (Foong et al., 2019). To find the noise parameter, a grid search over $10$ values on a log-scale between $21e-$ 2 and $21e$ 2 was used to find the noise that minimized the Gaussian likelihood of the validation set, with the LLA mean predictor as the Gaussian mean. The same grid search was used to find the prior variance, in order to minimize the expected calibration error (ECE) of LLA on the validation set. For the Protein dataset, a Kronecker-Factored Curvature (KFAC) covariance structure was used (Immer et al., 2021), and for the Song dataset a diagonal covariance structure was used. For all other datasets, LLA used the full covariance structure. The predictive distribution was computed with $1000$ MC samples.
@@
E.3 Image Classification
We display the training procedure in Table 9 for both Figure 2 and Table 2. For MNIST and FashionMNIST, we took a $5:1$ training/validation split of the training data. For CIFAR-10, we simply used the test data as a validation set. For CIFAR-10, random horizontal crop and flip on the training images was used as regularization.
@@ E.3 Image Classification
  • SWAG: The network was trained for a further 1×1$\times$1 × training epochs, at $1e2\times$learning rate. The covariance rank was set at $10$. At test time, $100$ MC-samples were taken.
  • LLA*: We used a last-layer KFAC approximation to the covariance. The prior precision was found through a grid search over $20$ values on a log scale from $21e-$ 2 to $21e$ 2, using the probit approximation to the predictive, and a validation set. This configuration for LLA is what is recommended in Daxberger et al. (2021). We used $1000$ samples, to remedy the large amount of approximations used.
@@
Table 9: Training procedure for image classification results in Table 2 and Figure 2.
MAP LeNet5 MNIST Lenet5 FMNIST ResNet9 FMNIST ResNet50 CIFAR10 ResNet50 CIFAR100
Learning Rate $35e-$ 3 $35e-$ 3 $31e-$ 3 $21e-$ 2 $11e-$ 1
Epochs $35$ $35$ $10$ $200$ $200$
Weight Decay $41e-$ 4 $41e-$ 4 $41e-$ 4 $45e-$ 4 $45e-$ 4
Batch Size $152$ $152$ $100$ $128$ $128$
Optimizer Adam Adam Adam SGD SGD
Scheduler Cosine Cosine Cosine Cosine Cosine
Accuracy $99\%$ $90\%$ $92.5\%$ $92.5\%$ $75\%$
@@
Learning Rate $21e-$ 2 $21e-$ 2 $11e-$ 1 $11e-$ 1 $21e-$ 2
Epochs $10$ $10$ $5$ $50$ $10$
Batch Size $152$ $152$ $50$ $128$ $128$
$\gamma$ $1$ $0.7$ $0.6$ $0.01$ $0.05$
E.4 Comparison Experiments
E.4.1 SLU
@@
For the results in Table 3, we copied the training details for (Miani et al., 2024, Table 3.); we point the reader to (Miani et al., 2024, Appendix D.1) for the exact details. For NUQLS, we used the following hyper-parameters for both MNIST and FashionMNIST: Epochs $10$, samples $10$, learning rate $0.01$, batch size $152$, $\bm{\gamma}$ $1$.
E.4.2 Sampling-LLA, VaLLA and ELLA
@@
For the results in Table 4, we used the training details for (Antorán et al., 2022, Figure 3. (left)); again, we point the reader to the details given in (Antorán et al., 2022, Appendix F.1). For NUQLS, we used the following hyper-parameters for MNIST: Epochs $10$, samples $10$, learning rate $0.01$, batch size $152$, $\bm{\gamma}$ $0.25$. For ResNet20 the following hyper-parameters were used: Epochs $1$, samples $50$, learning rate $0.0001$, batch size $152$, $\bm{\gamma}$ $0.01$.

### v2

@@ Appendix I Experiment Details
I.1 NTK Convergence
For this experiment, 2 randomly sampled sets of $100$ Gaussian inputs of dimension $5$, and $100$ scalar Gaussian targets, were used as training points and test points respectively. The weights of the MLP were initialised to $\mathcal{N}(0,1)$; no bias terms were used. To train, GD was employed with a Nesterov momentum parameter of $0.9$, a learning rate of $0.1$, and $5000$ epochs. For NUQLS, a learning rate of $0.1$ was used, and $\gamma$ was set to $1$. Error bars on Figure˜2 are the $95\%$ sample confidence interval.
I.2 Toy Regression
We use a 1-layer MLP, with a width of $50$ and SiLU activation. For the maximum a posteriori (MAP) network, we train for $10000$ epochs, with a learning rate of $0.001$, using the Adam optimizer and the PyTorch polynomial learning rate scheduler, with parameters total_\_iters == epochs, power =0.5=~0.5. For DE, each network in the ensemble outputs a heteroskedastic variance, and is trained using a Gaussian NLL, with $2000$ epochs and a learning rate of $0.05$. We combine the predictions of the ensembles as per (Lakshminarayanan et al., 2017). Both DE and NUQLS use $10$ realizations. The $\gamma$ hyper-parameter in NUQLS is set to $5$, and each linear realization is trained for $1000$ epochs with a learning rate of $0.001$, using SGD with a momentum parameter of $0.9$. In SWAG, the MAP network is trained for a further $10000$ epochs, using the same learning rate, and the covariance is formed with a rank-$10$ approximation. A prior precision of $0.1$ and $1$ is used for LLA and LA respectively, as well as the full covariance matrix. The variational inference method used is Bayes By Backprop (Blundell et al., 2015), as deployed in the Bayesian Torch package (Krishnan et al., 2022). The prior parameters are ($\mu=0,\sigma=1$), and the posterior is initialized at ($\mu=0,\rho=-3$). For SWAG, LLA, LA and VI, $1000$ MC sample were taken at test time. These design choices gave the best performance for this problem.
I.3 UCI Regression
Table 12: Training procedure for UCI regression results in Table˜1 and Table˜9.
@@ I.3 UCI Regression
No. experiments $10$ $10$ $10$ $10$ $10$ $10$ $10$ $10$ 33
We now provide the experimental details for the UCI regression experiments (as seen in Table˜1 and Table˜9). For each dataset, we ran a number of experiments to get a mean and standard deviation for performance metrics. In each experiment, we took a random $70\%/15\%/15\%$ split of the dataset for training, testing, and validation. The training hyper-parameters for the MAP, DE and NUQLS networks, size of the MLP used, and the number of experiments conducted for each dataset can be found in Table˜12. Reported time for NUQLS, LLA and SWAG includes the training time for the original NN, with the time to run the method given in brackets.
[…]
@@
  • LLA: LLA requires two parameters for regression: a dataset noise parameter, and a prior variance on the parameters (Foong et al., 2019). To find the noise parameter, a grid search over $10$ values on a log-scale between 1e21e-2 and $1e2$ was used to find the noise that minimized the Gaussian likelihood of the validation set, with the LLA mean predictor as the Gaussian mean. The same grid search was used to find the prior variance, in order to minimize the expected calibration error (ECE) of LLA on the validation set. For the Protein dataset, a Kronecker-Factored Curvature (KFAC) covariance structure was used (Immer et al., 2021), and for the Song dataset a diagonal covariance structure was used. For all other datasets, LLA used the full covariance structure. The predictive distribution was computed with $1000$ MC samples.
@@
I.4 Image Classification
We display the training procedure in Table˜13 for both Figure˜3 and Table˜2. For MNIST and FashionMNIST, we took a $5:1$ training/validation split of the training data. For CIFAR-10, we simply used the test data as a validation set. For CIFAR-10, random horizontal crop and flip on the training images was used as regularization.
@@ I.4 Image Classification
  • SWAG: The network was trained for a further $1\times$ training epochs. For CIFAR-$100$ and SVHN, learning rate was kept the same as the original NN; for other datasets, a larger learning rate of $1e2\timesNN$ learning rate was used. The covariance rank was set at $10$. At test time, $100$ MC-samples were taken.
  • LLA*: We used a last-layer KFAC approximation to the covariance. The prior precision was found through a grid search over $20$ values on a log scale from 1e21e-2 to $1e2$, using the probit approximation to the predictive, and a validation set. This configuration for LLA is what is recommended in Daxberger et al. (2021). We used $1000$ samples, to remedy the large amount of approximations used.
@@
Table 13: Training procedure for image classification results in Table˜2, Figure˜3, and Figure˜6.
MAP LeNet5 MNIST Lenet5 FMNIST ResNet9 FMNIST ResNet50 CIFAR10 ResNet50 CIFAR100 ResNet50 SVHN ResNet50 ImageNet
Learning Rate $5\times 10^{-3}$ $5\times 10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ $10^{-2}$ -
Epochs $35$ $35$ $10$ $200$ $200$ $50$ -
Weight Decay $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ -
Batch Size $152$ $152$ $100$ $128$ $128$ $128$ -
Optimizer Adam Adam Adam SGD SGD SGD -
Scheduler Cosine Cosine Cosine Cosine Cosine Cosine -
Accuracy $99\%$ $90\%$ $\% 92.592.5\%$ $75\%$ $87\%$ $76\%$
@@
Learning Rate $10^{-2}$ $10^{-2}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-2}$
Epochs $10$ $10$ 22 $50$ $10$ 22 22
Batch Size $152$ $152$ $50$ $128$ $128$ $256$ $56$
$\gamma$ 11 $0.7$ $0.1$ $0.01$ $0.05$ $0.05$ $0.7$
I.5 Confidence Intervals
For the confidence intervals experiments in Table˜4 and Table˜3, we used $100$ repeats for $d\leq 4$ and $50$ repeats for $d\geq 8$. We used a multiplier of $1.5$ on all NUQLS variance values, as we note that the search strategy for $\gamma$ for regression only employs ECE; this does not reward confidence intervals that exceed the expected confidence, and hence computed confidence intervals may be conservative in width.
I.6 Comparison Experiments
I.6.1 BDE
For Figure˜4 and Table˜8, BDE was implimented following details given in (He et al., 2020). To make comparison with NUQLS fair, each ensemble member of BDE did not output a separate hetero-skedastic variance term. Instead, the variance over predictions was taken as the variance. Further, the large computational cost of BDE relative to NUQLS meant that it is unrealistic for comparison to allow BDE to tune hyper-parameters through a grid-search, as this would take the run time of BDE (which is slightly greater than that of DE) severely out-of-proportion to NUQLS. For Figure˜4, the noise-parameter was given as the noise in the data, $9$. For Table˜8, standard Gaussian noise was taken a priori. The ensemble members were trained with the same training scheme as the original NN in both Figure˜4 and Table˜8. For both experiments, $10$ ensemble members were used. For classification, we again followed the implementation details given in He et al. (2020) to the best of our ability. We employed the same hyperparameters as DE/MAP, and used $10$ ensemble members.
I.6.2 SLU
@@
For the results in Table˜5, we copied the training details for (Miani et al., 2024a, Table 3.); we point the reader to (Miani et al., 2024a, Appendix D.1) for the exact details. For NUQLS, we used the following hyper-parameters for both MNIST and FashionMNIST: Epochs $10$, samples $10$, learning rate $0.01$, batch size $152$, $\bm{\gamma}$ 11.
I.6.3 Sampling-LLA, VaLLA and ELLA
@@
For the results in Table˜6, we used the training details for (Antorán et al., 2022, Figure 3. (left)); again, we point the reader to the details given in (Antorán et al., 2022, Appendix F.1). For NUQLS, we used the following hyper-parameters for MNIST: Epochs $10$, samples $10$, learning rate $0.01$, batch size $152$, $\bm{\gamma}$ $0.25$. For ResNet20 the following hyper-parameters were used: Epochs $1$, samples $100$, learning rate $0.0001$, batch size $152$, $\bm{\gamma}$ $0.01$. For ResNet32 the following hyper-parameters were used: Epochs $1$, samples $150$, learning rate $0.0001$, batch size $152$, $\bm{\gamma}$ $0.01$.
I.6.4 BE
We employed the WideResNet architecture from Franchi et al. (2023). We used a depth of $10$ and a widening factor of $1$. The MAP network was trained with: Epochs $20$, learning rate $0.005$, weight decay $0.0001$, batch size $152$, the Adam optimizer and the Cosine learning rate scheduler. BE employed the same hyperparameters, yet with $1.5\timesEpochs$, to aid with training. BE used $5$ ensemble members, in order to match the computational budget of NUQLS. The NUQLS hyperparameters were: Epochs $10$, samples $10$, learning rate $0.01$, batch size $152$, $\bm{\gamma}$ 11.
I.6.5 SNGP
We employed the Lightning UQ Box (Lehmann et al., 2025) implementation of SNGP for our experiments. The same hyperparameters as the MAP network were used. Note that we also tested SNGP on ResNet50 SVHN, however SNGP was unable to train successfully, even after trying $3$ learning rates and $2$ optimizers. Hence, we did not include the results for ResNet50 SVHN.
I.6.6 SGLD
We have implemented the SGLD method from Welling & Teh (2011) using the Lightning UQ Box package (Lehmann et al., 2025) as the basis of the code. We then amended this code to include the learning rate scheduler from Welling & Teh (2011). We followed Maddox et al. (2019), and initialized the SGLD trajectory from the weights of the trained network. We also copied the learning rate from Maddox et al. (2019), and used the same weight-decay as the original network for the SGLD prior. However, in contrast to Maddox et al. (2019), we took the noise-factor scaling to be $10^{-3}$ instead of $5\times 10^{-4}$, as we found that this gave better performance. We sampled $100$ epochs from the posterior, using a batch size of $100$.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We claim that our method approximates the posterior of an NTK-GP, and provide theoretical and empirical proof in Section˜3 and Section˜4.1 respectively. We also claim that we outperform competing methods in performance and efficiency, which can be seen in Section˜4 and Appendix˜H.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Discussion of our limitations can be found in the "Limitations" subsection of the Conclusion, Section˜5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We contain two provable results, Lemma˜3.1 and Theorem˜3.2. The proofs can be found in Appendix˜A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All code for the experiments has been included with the submission. All experimental details have been described in Appendix˜I.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. 1.
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. 2.
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. 3.
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. 4.
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Link to code, as well as a downloadable package of our method, can be found at the top of Section˜4.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All training and test details are found in Appendix˜I.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: UCI Regression experiments (Tables˜1, 9 and 8 and the empirical convergence figure (Figure˜2) have error bars included, and are defined in Figure˜2 and in Section˜I.3. Violin plots do not have error bars, as they are generally not reported with error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Experimental resources are listed at the top of Appendix˜I.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: After reading through the NeurIPS Code of Ethics, we believe that we have conformed to this code.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the benefits of improved UQ methods in our Introduction and Abstract. We are not aware of any direct negative impacts of our work (though secondary negative impacts may exist).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We did not use any high risk models or datasets, and instead only use highly common models and datasets that are commonly used in the UQ literature.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Any prior code for related methods that has been used has been referenced in the linked GitHub repository for our code. Any related methods for comparison has been properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our method has been released as package, with examples given for how to correctly use our method.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The development of our method did not involve the use of LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.